OpenStudy (ilovebmth1234):

Solve for x. with steps. first person to help me gets a medal. :) -3(x – 7) = 2(x + 3)

OpenStudy (zpupster):

first multiply the term -3 by each term in the parenthesis

OpenStudy (zpupster):

on left hand side

OpenStudy (ilovebmth1234):

huh??

OpenStudy (zpupster):

ok i will do the left hand side -3x+21= now mult each term by 2 on the rt hand side

OpenStudy (ilovebmth1234):

wow im confused as flutter

OpenStudy (zpupster):

we are doing this a(b+c)= ab+ac you see it is a multiplication problem we multiply a by each term inside the parenthesis now try 2(x + 3)

OpenStudy (anonymous):

-3(x – 7) = 2(x + 3) distribute -3x+21=2x+6 then put like terms together 15=5x divide by -1 x=3

OpenStudy (anonymous):

i meant divide by 5

OpenStudy (ilovebmth1234):

how did you get 21 and 6?

OpenStudy (anonymous):

when you distribute the -3 and 2